OpenStudy (jcarter91198):

y-5=3/10(x--5) into standard form that passes through the points (-4, 5) and (6, 2)

OpenStudy (sshayer):

slope of this line=3/10 slope of line passing through (-4,5)and (6,2) \[=\frac{ 2-5 }{ 6-(-4) }=\frac{ -3 }{ 10 }\] so two points can not lie on the given line. They can cut only in one point.

OpenStudy (sshayer):

10 y-50=3x-15 10y-3x=35 (1) eq. of line through two points is \[y-5=\frac{ -3 }{ 10 }\left( x-(-4) \right)=\frac{ -3 }{ 10 }\left( x+4 \right)\] 10y-50=-3x-12 10 y+3x=38 (2) adding (1) and (2) 20 y=73 y=3.65 3x=38-36.5=1.5 x=0.5 so these given point does not lie on the given line.

OpenStudy (sshayer):

i think the points should be (4,5)and (6,2)